Theresa invests $2000 into a GIC which will increase in value at a rate of 4% compounded annually, wheras Steve invests $1000 into a GIC which increases in value at a rate of 6% compounded annually. How many years will pass before Theresa and Steve have the same amount of money ?
Answer by scott8148(6628)

let x="years to equality"

2000(1.04)^x=1000(1.06)^x __ dividing by 1000 __ 2(1.04)^x=(1.06)^x

taking log __ log(2)+x(log(1.04))=x(log(1.06))

subtracting x(log(1.04)) __ log(2)=x(log(1.06))-x(log(1.04))

factoring __ log(2)=x(log(1.06)-log(1.04))

dividing by (log(1.06)-log(1.04)) __ (log(2))/(log(1.06)-log(1.04))=x